OpenStudy (anonymous):

write the quadrict function in standard form y=(x+1)^2-8

OpenStudy (anonymous):

The standard form is y=ax^2+bx+x

OpenStudy (anonymous):

Try to take it to that form

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

nope still dont get it

OpenStudy (majikdusty):

Hint: You must foil (x+1)^2

OpenStudy (anonymous):

how!!!!!!!

OpenStudy (majikdusty):

So (x+1)^2 is the same thing as (x+1)(x+1) FOIL stands for first-outer-inner-last first- x*x=x^2 outer x*1=x inner x*1=x last 1*1=1 and you cannot forget the minus 8. So at this point you just have to combine like terms

OpenStudy (anonymous):

so i still dont get it ...

OpenStudy (majikdusty):

well the final answer is (X^2)-2x-7. Do you not understand how to foil? or is it something else?